OpenStudy (anonymous):

write the slope intercept equation for a line that passes through (-4,-8)(-3,-2)

OpenStudy (anonymous):

\[y=6x+16\]

OpenStudy (amistre64):

the definition of slope is; the change in "y" over the change in "x". y1-y2 ----- = slope x1-x2 -8-(-2) -6 ------ = --- -4-(-3) -1 slope= --6 = +6